Class drill3 Cardella Company applies overhead on the basis of direct labor costs. The company estimates annual overhead costs will be $760,000 and annual direct labor costs will be $950,000. During February, Cardella works on two jobs: A16 and B17. Summary data concerning these jobs are as follows. Manufacturing Costs Incurred Purchased $54,000 of raw materials on account. Factory labor $76,000, plus $4,000 employer payroll taxes. Manufacturing overhead exclusive of indirect materials and indirect labor $59,800. Assignment of Costs Direct materials: Job A16 $27,000, Job B17 $21,000 Indirect materials: $3,000 Direct labor: Job A16 $52,000, Job B17 $26,000 Indirect labor: $2,000 . . The company completed Job A16 and sold it on account for $150,000. Job B17 was only partially completed. INSTRUCTIONS 1. Compute the predetermined overhead rate. 2. Journalize the February transactions in the sequence 3. What was the amount of under or overapplied manufacturing overhead?
Answer #1
1) Pre-determined overhead rate = 760000/950000 = 80.00%
2) Raw materials inventory 54000
Accounts payable 54000
[To record purchase of raw materials on account]
Factory payroll 80000
Salaries and wages payable 76000
Payroll taxes payable 4000
[To record incurring of factory salaries and wages]
Manufacturing overhead 59800
Accounts payable 59800
[To record mfg overhead incurred]
Work in process inventory 48000
Manufacturing overhead 3000
Raw materials inventory 51000
[To record issue of raw materials]
Work in process inventory 78000
Manufacturing overhead 2000
Factory payroll 80000
[To assign labor costs]
Work in process inventory [(52000+26000)*80%] 62400
Manufacturing overhead 62400
[To record assignment of manufacturing overhead]
Finished goods inventory [27000+52000+52000*80%] 120600
Work in process inventory 120600
[To record cost of Job A16 on completion]
Accounts receivable 150000
Sales 150000
[To record sale of Job A16 on account]
Cost of goods sold 120600
Finished goods inventory 120600
[To record COGS on sale]
3) Overhead incurred = 59800+3000+2000 = 64800
Overhead applied 62400
Underapplied MOH 2400